OpenStudy (anonymous):

The temperature of a city changed by −15 degrees Celsius over a 5- week period. The temperature changed by the same amount each week. Part A: Write an expression to show the average change in the temperature of the city each week. Part B: Simplify the expression and explain, using words, what your answer means. @Princer_Jones

OpenStudy (anonymous):

Is it -15 degrees? I mean negative 15?

OpenStudy (anonymous):

yea :)

OpenStudy (anonymous):

Ok in 5 weeks, the temperature change by -15 degree Celsius, also it is given that it decreased by the same amount each week. Hence each week it decreased by -15/5=-3 degree Celsius

OpenStudy (anonymous):

ok, so whats the expression for part a?

OpenStudy (anonymous):

Part A answer, The average change in temperature of the city each week is x/5, where x is the total change in temperature in 5 weeks

OpenStudy (anonymous):

ok and part b? :)

OpenStudy (anonymous):

Part B : Putting x=-15 we have, average change in temperature per week=-15/5=-3 degree celcius which implies that it is the season of winter because the temperature is decreasing

OpenStudy (anonymous):

and what does the asnwer mean?

OpenStudy (anonymous):

Yeah i said it in Part B

OpenStudy (anonymous):

oh ok thank you again!!! I just have have one more quick one? You think you can help me? :)

OpenStudy (anonymous):

ok but last one..:)

OpenStudy (anonymous):

yay thanks :)